An infinite geometric sequence has 1 and 1/5 as its first two terms: 1 1/5 1/25 1/125..... what is the sum,s, of the infinite series?
Answer by oscargut(2103)

the serie k^n from n=0 to infinite converges if
-1 < k < 1 and the sum value is 1/(1-k)
Then the sum when k=1/5 is: 1/(1-1/5)=1/(4/5)=5/4
Answer: 5/4